a) Express the following surfaces in spherical coordinates
i) xz=3
ii) x2 + y2 - z2=1
b) Find the cylindrical coordinates of the points where the Cartesian coordinates are
i). (6,6,8)
ii) (√2,1,1)
1
Expert's answer
2020-06-22T17:24:38-0400

a) Interrelation for Cartesian and spherical coordinates:

"x = \rho\sin\theta\cos\phi\\
y = \rho\sin\theta\sin\phi\\
z = \rho\cos\theta"


i) xz = 3 =>

"\rho\sin\theta\cos\phi*\rho\cos\theta=3" =>

"\rho^2\sin(2\theta)\cos\phi = 6"


ii) "x^2 + y^2 - z^2=1" =>

"(\rho\sin\theta\cos\phi)^2 + (\rho\sin\theta\sin\phi)^2 - (\rho\cos\theta)^2 = 1" =>

"\rho^2(\sin^2\theta(\cos^2\phi + \sin^2\phi) - \cos^2\theta) = 1" =>

"\rho^2(\sin^2\theta - \cos^2\theta) = 1" =>

"\rho^2\cos(2\theta) = -1"



b) a) Interrelation for cylindrical and Cartesian coordinates:

"r = \sqrt{x^2+y^2}\\
\theta = \arctan{\cfrac{y}{x}}\\
z = z"


i) "(6,6,8)_{Cartesian}" =>

"(\sqrt{6^2+6^2}, \arctan{\cfrac66, 8})_{cylindrical}" =>

"(6\sqrt2, 45^\circ, 8)"


ii) "(\sqrt2, 1, 1)_{Cartesian}" =>

"(\sqrt{(\sqrt2)^2+1^2}, \arctan{\cfrac1{\sqrt2}}, 1)_{cylindrical}" =>

"(\sqrt3, 35.264^\circ, 1)"